14. Your target is to have $200,000 for your daughter’s college tuition. If you begin saving today equal amounts per year for 15 years at 8%, how much will you need to save per year?
Answer #1

Ans $ 7365.91

P = Periodic payments
r = rate of interest
n = no of years
Future Value of Annuity = P ( (1 + r)n - 1 ) / r
200000 = P* ((1 + 8%)^15 - 1) / (8%)
200000 = 27.1521139274784 * P
P = 200000 / 27.1521139274784
P = 7365.91